For the function, y=1/x-2
a) Give the y values for x = -2, -1, 0, 1, 2, 3.
Answer:
Show work in this space.

Answer by CrazyMan Jr.(21)

y=x-2
y=(-2)-2 = (-4)
y=(-1)-2 = (-3)
y=(0)-2 = (-2)
y=(1)-2 = (-1)
y=(2)-2 = (0)
y=(3)-2 = (1)
J-Man